# Simultaneous symmetry breaking in spontaneous Floquet states: Floquet-Nambu-Goldstone modes, Floquet thermodynamics, and the time operator

### v3

@@ III.2 Nambu-Goldstone modes in the generalized Gibbs ensemble
$(z_{a}|z_{n})$ $=$ $0$
III.3 Berry-Gibbs connection
The derivation of the orthogonality relations within the Goldstone-Gibbs sector is more delicate. First, we note that the modes $z_{a}$ are self-conjugates, $z_{a}=\bar{z}_{a}$, so their inner product is purely imaginary, $(z_{a}|z_{b})=-(z_{a}|z_{b})^{*}$, and they have zero norm, $(z_{a}|z_{a})=0$. In particular, for the NG modes, since by assumption $[T_{\alpha},T_{\beta}]=0$, $\braket{\Psi_{0}}{[T_{\alpha},T_{\beta}]}{\Psi_{0}}=0$, and then
$(z_{\alpha}|z_{\beta})=\braket{T_{\alpha}\Psi_{0}}{T_{\beta}\Psi_{0}}-\braket{T_{\beta}\Psi_{0}}{T_{\alpha}\Psi_{0}}=0$ (47)
@@ III.3 Berry-Gibbs connection
$\partial_{b}Q_{\alpha}=\int\mathrm{d}\mathbf{x}~{}(T_{\alpha}\Psi_{0})^{*}\partial_{b}\Psi_{0}+(\partial_{b}\Psi_{0})^{*}T_{\alpha}\Psi_{0}=-i(z_{\alpha}|z_{b}).$ (48)
@@
$\partial_{A}\Psi_{0}$$\to$ $e^{-i\boldsymbol{\chi}\cdot\mathbf{T}}\left[\partial_{A}-i\partial_{A}\chi^{\beta}T_{\beta}\right]\Psi_{0},$
$(z_{A}|z_{B})$$\to$ $(z_{A}|z_{B})+i(\partial_{A}\chi^{\beta}-\partial_{B}\chi^{\alpha}),$ (50)
@@
$F_{AB}=i(z_{A}|z_{B})=i(\braket{\partial_{A}\Psi_{0}}{\partial_{B}\Psi_{0}}-\braket{\partial_{B}\Psi_{0}}{\partial_{A}\Psi_{0}})$ (53)
is nothing else than the usual Berry curvature evaluated for the GPG wavefunction, which here depends on the set of conserved charges $\mathbf{Q}$ (we recall that they are not intrinsic parameters of the original Hamiltonian). We then immediately identify $F=dW$, with $W_{A}=i\braket{\Psi_{0}}{\partial_{A}\Psi_{0}}$  [^2] Strictly speaking, it should be $W_{A}=i(\braket{\Psi_{0}}{\partial_{A}\Psi_{0}}-\braket{\partial_{A}\Psi_{0}}{\Psi_{0}})/2$ to ensure that $W_{A}$ is real since $\partial_{A}\braket{\Psi_{0}}{\Psi_{0}}=\partial_{A}N\neq 0$.. As a result of this analogy, we denote the vector $W$ as the Berry-Gibbs connection and $F_{AB}$ as the Berry-Gibbs curvature. Its main differences with the standard Berry connection are i) $\braket{\Psi_{0}}{\Psi_{0}}$ does depend on the connection parameters (specifically, $\braket{\Psi_{0}}{\Psi_{0}}=N$); and ii) we can perform generalized gauge transformations, which go beyond the usual phase transformations $\Psi_{0}\to e^{-i\theta(\mathbf{Q})}\Psi_{0}$, under which $F_{AB}$ is not invariant.
@@
$F_{ab}\equiv i(z_{a}|z_{b})=i(\braket{\partial_{a}\Psi_{0}}{\partial_{b}\Psi_{0}}-\braket{\partial_{b}\Psi_{0}}{\partial_{a}\Psi_{0}}),$ (54)
@@
$x^{\prime\alpha}=x^{\alpha}+\chi^{\alpha}(\mathbf{Q}),~{}x^{\prime A}=x^{A}.$ (56)
@@
$x_{b}=x^{a}\Omega_{ab},~{}x^{a}=x_{b}\Omega^{ba},$ (59)
@@
III.4 Quantization
When returning to the quantum theory, quantization à la Dirac prescribes that Poisson brackets are promoted to commutators as $[\hat{A},\hat{B}]=i\hbar\{A,B\}$. Therefore, the field and the conserved charges are promoted to quantum operators which satisfy [see Eqs. (11), (16)]
@@ III.4 Quantization
In the case of a fermionic field operator, the commutator is replaced by an anticommutator. The symmetry transformations on the field operator can be implemented as
@@
where $\mathbf{\hat{Q}}$ is a vector containing the second-quantized charges $\hat{Q}_{\alpha}$, in turn satisfying $[\hat{H},\hat{Q}_{\alpha}]=[\hat{Q}_{\alpha},\hat{Q}_{\beta}]=0$. The Heisenberg equation of motion for the field operator (III) results in the quantum version of the time-dependent GP equation (22):
$i\partial_{t}\hat{\Psi}(\mathbf{x},t)=\left[-\frac{\nabla^{2}}{2}+V(\mathbf{x})+\hat{\Psi}^{\dagger}(\mathbf{x},t)\hat{\Psi}(\mathbf{x},t)\right]\hat{\Psi}(\mathbf{x},t).$ (62)
We now adapt the reasoning behind Eq. (40) by performing a unitary transformation
@@
equivalent to working in a Gibbs picture where the dynamics is governed by the generalized Gibbs Hamiltonian, $\hat{\Psi}^{\prime}(\mathbf{x},t)=e^{i\hat{K}t}\hat{\Psi}^{\prime}(\mathbf{x})e^{-i\hat{K}t}$. The resulting equation of motion in the Gibbs picture is the quantum version of the time-dependent GPG equation:
$i\hbar\partial_{t}\hat{\Psi}^{\prime}(\mathbf{x},t)=[\hat{\Psi}^{\prime}(\mathbf{x},t),\hat{K}]$ (64)
$=\left[-\frac{\nabla^{2}}{2}+V(\mathbf{x})+\hat{\Psi}^{\prime\dagger}(\mathbf{x},t)\hat{\Psi}^{\prime}(\mathbf{x},t)-\boldsymbol{\lambda}\cdot\mathbf{T}\right]\hat{\Psi}^{\prime}(\mathbf{x},t).$
These results are general and exact, easily translatable to an arbitrary many-body system. In the specific case of a condensate, we can expand the field operator around the GPG expectation value, $\hat{\Psi}^{\prime}(\mathbf{x},t)=\Psi_{0}(\mathbf{x})+\hat{\varphi}(\mathbf{x},t)$, recovering at linear order the quantum time-dependent BdGG equations (42),
$i\partial_{t}\hat{\Phi}=M_{0}\hat{\Phi},~{}\hat{\Phi}(\mathbf{x},t)=\left[\begin{array}[]{c}\hat{\varphi}(\mathbf{x},t)\\ \hat{\varphi}^{\dagger}(\mathbf{x},t)\end{array}\right].$ (65)
@@
$[\hat{\gamma}_{n},\hat{\gamma}^{\dagger}_{m}]=[(z_{n}|\hat{\Phi}),(\hat{\Phi}|z_{m})]=(z_{n}|z_{m})=\delta_{nm}.$ (68)
@@
$i\partial_{t}\hat{\gamma}_{n}(t)=(z_{n}|M_{0}\hat{\Phi})=\omega_{n}\hat{\gamma}_{n}(t)\Longrightarrow\hat{\gamma}_{n}(t)=\hat{\gamma}_{n}e^{-i\omega_{n}t}.$ (69)
@@
$[\hat{\gamma}^{a},\hat{\gamma}^{b}]=-i\Omega^{ab}.$ (71)
This implies that the components of the vector $\boldsymbol{\hat{\gamma}}=(\hat{\gamma}^{\alpha},\hat{\gamma}^{A})\equiv(\hat{X}^{\alpha},\hat{P}^{A})$ can be seen as generalized coordinate and momenta operators, as those also satisfy the same symplectic algebra. Indeed, we can take this analogy further by analyzing their equations of motion,
$i\partial_{t}\hat{\gamma}_{a}(t)=i(z_{a}|M_{0}\hat{\Phi})=i(M_{0}z_{a}|\hat{\Phi})$ (72)
@@
$\hat{X}^{\alpha}$$\equiv$ $\hat{\gamma}^{\alpha}=i(z_{A}|\hat{\Phi}),$ (73)
$\hat{P}^{A}$$\equiv$ $\hat{\gamma}^{A}=-i(z_{\alpha}|\hat{\Phi}),$
@@
The quantum fluctuations of the velocities are then translated into a ballistic motion for the quantum amplitude of the NG modes, arriving at Eq. (75). Another way to put it is that the modes $z_{A}$ are not eigenmodes of the BdGG equations. Instead, from Eq. (24), we have that
@@
which yields $z_{A}(\mathbf{x},t)=z_{A}(\mathbf{x})+(\partial_{A}\lambda^{\beta})z_{\beta}t$ in the time-dependent BdGG equation (42). As a result, we recover the same intuitive picture of the classical mechanical example in Sec. II: the dynamics of the fluctuations along the orbits generated by the symmetry transformations lead to unbounded motion as there is no restoring force.
After gathering Eqs. (69), (74), we find that the quantum field fluctuations evolve as
@@
$+$ $\sum_{n}\hat{\gamma}_{n}z_{n}(\mathbf{x})e^{-i\omega_{n}t}+\hat{\gamma}^{\dagger}_{n}(t)\bar{z}_{n}(\mathbf{x})e^{i\omega_{n}t}.$
@@
The same dynamics could have been derived from the generalized Gibbs Hamiltonian by inserting Eq. (66) in $\hat{K}=\hat{H}-\lambda^{\alpha}\hat{Q}_{\alpha}$ and expanding up to quadratic order in the field fluctuations, in the spirit of the Bogoliubov approximation. After dropping out all mean-field and zero-point $c$-number contributions, we arrive at
$\hat{K}=\frac{1}{2}(\hat{\Phi}|M_{0}\hat{\Phi})=\frac{\partial^{2}_{AB}E}{2}\hat{P}^{A}\hat{P}^{B}+\sum_{n}\omega_{n}\hat{\gamma}^{\dagger}_{n}\hat{\gamma}_{n}.$ (80)
@@
The second term is the usual Bogoliubov contribution while the first one is the Goldstone-Gibbs contribution. The latter is quadratic in the fluctuations of the conserved charges, then behaving as effective momenta, and does not depend on the NG amplitudes, since the Hamiltonian is invariant under the symmetry transformations. In particular, if the solution is energetically stable, then $\partial^{2}_{AB}E$ is a positive definite form. In fact, one can always diagonalize $\partial^{2}_{AB}E$ by performing an appropriate rotation in the Goldstone-Gibbs sector, $\partial^{2}_{AB}E=M_{A}^{-1}\delta_{AB}$, where $M_{A}$ plays the role of an effective mass. This rotation yields a new set of independent Goldstone-Gibbs modes $\{z^{\prime}_{\alpha},z^{\prime}_{A}\}$, whose amplitudes $\{\hat{X}^{\prime\alpha},\hat{P}^{\prime A}\}$ satisfy
@@
$\hat{K}=\frac{(\hat{P}^{\prime A})^{2}}{2M_{A}}+\sum_{n}\omega_{n}\hat{\gamma}^{\dagger}_{n}\hat{\gamma}_{n},~{}\hat{X}^{\prime\alpha}(t)=\hat{X}^{\prime\alpha}+\frac{\hat{P}^{\prime A}}{M_{A}}t.$ (81)
@@
Thus, the matrix $\partial^{2}_{AB}E$, computed from thermodynamical considerations, determines how the different NG modes hybridize in a dynamical context. Moreover, as the NG modes are the zero-frequency limit of some branches from the regular Bogoliubov sector, the structure of the hybrid Goldstone-Gibbs modes $\{z^{\prime}_{\alpha},z^{\prime}_{A}\}$ also describes their low-frequency regime.
@@
The quantization of the Goldstone-Gibbs modes in the usual GP context is simply retrieved by first setting all Lagrange multipliers to zero except for the chemical potential $\lambda^{\theta}=\mu$, so the only independent conserved charge is the particle number $N$. With the resulting GP solution, one computes the remaining charges $Q_{\alpha}$ to assemble the vector $\mathbf{Q}=\mathbf{Q}_{0}$. Once known, the full GPG equation is solved as a function of $\mathbf{Q}$, from where the Berry-Gibbs curvature $F_{AB}$ and the matrix $\partial^{2}_{AB}E$ can be evaluated at the point $\mathbf{Q}_{0}$ characterizing the GP solution.
@@
To conclude this section, we stress that the thermodynamical description is essential to ensure the conservation of the commutation rules for the NG amplitudes because it implies that the flow of the GPG wavefunction is potential, i.e., $\dot{\alpha}=\lambda_{A}=\partial_{A}E$, and hence
$[\hat{X}_{A}(t),\hat{X}_{B}(t)]=i(\partial_{B}\lambda_{A}-\partial_{A}\lambda_{B})t=0.$ (82)
IV Cnoidal wave in a superfluid

### v4

@@ 3.2 Nambu-Goldstone modes in the generalized Gibbs ensemble
$(z_{a}|z_{n})$ $=$ $0.$
3.3 Berry-Gibbs connection
The derivation of the orthogonality relations within the Goldstone-Gibbs sector is more delicate. First, we note that the modes $z_{a}$ are self-conjugates, $z_{a}=\bar{z}_{a}$, so their inner product is purely imaginary, $(z_{a}|z_{b})=-(z_{a}|z_{b})^{*}$, and they have zero norm, $(z_{a}|z_{a})=0$. In particular, for the NG modes, since by assumption $T_{\alpha},T_{\beta}]=0$, $\braket{\Psi_{0}|[T_{\alpha},T_{\beta}]|\Psi_{0}}=0$, and then
$(z_{\alpha}|z_{\beta})=\braket{T_{\alpha}\Psi_{0}|T_{\beta}\Psi_{0}}-\braket{T_{\beta}\Psi_{0}|T_{\alpha}\Psi_{0}}=0$ (50)
@@ 3.3 Berry-Gibbs connection
$\partial_{b}Q_{\alpha}$ $=$ $\int\mathrm{d}\mathbf{x}\leavevmode\nobreak\ (T_{\alpha}\Psi_{0})^{*}\partial_{b}\Psi_{0}+(\partial_{b}\Psi_{0})^{*}T_{\alpha}\Psi_{0}$ (51)
$=$ $-i(z_{\alpha}|z_{b}).$
@@
$\partial_{A}\Psi_{0}$ $\to e^{-i\boldsymbol{\chi}\cdot\mathbf{T}}\left[\partial_{A}-i\partial_{A}\chi^{\beta}T_{\beta}\right]\Psi_{0},$ (53)
$(z_{A}|z_{B})$ $\to(z_{A}|z_{B})+i(\partial_{A}\chi^{\beta}-\partial_{B}\chi^{\alpha}),$
@@
$F_{AB}=i(\braket{\partial_{A}\Psi_{0}|\partial_{B}\Psi_{0}}-\braket{\partial_{B}\Psi_{0}|\partial_{A}\Psi_{0}})$ (56)
is nothing else but the usual Berry curvature evaluated for the GPG wavefunction, which here depends on the set of conserved charges $\mathbf{Q}$ (we recall that these are not intrinsic parameters of the original Hamiltonian). We then immediately identify $F=dW$, with $W_{A}=i\braket{\Psi_{0}|\partial_{A}\Psi_{0}}$ [^2] Strictly speaking, it should be $W_{A}=i(\braket{\Psi_{0}|\partial_{A}\Psi_{0}}-\braket{\partial_{A}\Psi_{0}|\Psi_{0}})/2$ to ensure that $W_{A}$ is real since $\partial_{A}\braket{\Psi_{0}|\Psi_{0}}=\partial_{A}N\neq 0$.. As a result of this analogy, we denote the vector $W$ as the Berry-Gibbs connection and $F_{AB}$ as the Berry-Gibbs curvature. Its main differences with the standard Berry connection are i) $\braket{\Psi_{0}|\Psi_{0}}$ does depend on the connection parameters (specifically, $\braket{\Psi_{0}|\Psi_{0}}=N$); and ii) we can perform generalized gauge transformations beyond the usual phase transformations $\Psi_{0}\to e^{-i\theta(\mathbf{Q})}\Psi_{0}$, under which $F_{AB}$ is not invariant.
@@
$F_{ab}\equiv i(z_{a}|z_{b})=i(\braket{\partial_{a}\Psi_{0}|\partial_{b}\Psi_{0}}-\braket{\partial_{b}\Psi_{0}|\partial_{a}\Psi_{0}}),$ (57)
@@
$x^{\prime\alpha}=x^{\alpha}+\chi^{\alpha}(\mathbf{Q}),\leavevmode\nobreak\ x^{\prime A}=x^{A}=Q_{\alpha}.$ (59)
@@
$x_{b}=x^{a}\Omega_{ab},\leavevmode\nobreak\ x^{a}=x_{b}\Omega^{ba},$ (62)
@@
3.4 Quantization
When returning to the quantum theory, quantization à la Dirac prescribes that Poisson brackets are promoted to commutators as $\hat{A},\hat{B}]=i\hbar\{A,B\}$. Therefore, the field and the conserved charges are promoted to quantum operators which satisfy [see Eqs. (11), (15)]
@@ 3.4 Quantization
In the case of a fermionic field operator, the commutator in the first line is replaced by an anticommutator. Symmetry transformations on the field operator can be implemented as
@@
where $\mathbf{\hat{Q}}$ is a vector containing the second-quantized charges $\hat{Q}_{\alpha}$, in turn satisfying $\hat{H},\hat{Q}_{\alpha}]=[\hat{Q}_{\alpha},\hat{Q}_{\beta}]=0$. The equation of motion for the quantum field operator is precisely the Heisenberg equation (3).
The Lagrangian description of Eq. (19) is retrieved by performing a unitary transformation
@@
equivalent to working in a Gibbs picture where the dynamics is governed by the generalized Gibbs Hamiltonian, $\hat{\Psi}^{\prime}(\mathbf{x},t)=e^{i\hat{K}t}\hat{\Psi}(\mathbf{x})e^{-i\hat{K}t}$. The point of the Gibbs picture is that it amounts to switching to a rotating frame where the symmetry-breaking state is time-independent, and the dynamics of its quantum fluctuations is determined by a stationary problem.
These results are so far general and exact, easily translatable to an arbitrary many-body system. In the specific case of a condensate, the Heisenberg equation of motion is the quantum version of the time-dependent GP equation (21),
$i\partial_{t}\hat{\Psi}(\mathbf{x},t)=\left[-\frac{\nabla^{2}}{2}+V(\mathbf{x})+\hat{\Psi}^{\dagger}(\mathbf{x},t)\hat{\Psi}(\mathbf{x},t)\right]\hat{\Psi}(\mathbf{x},t),$ (66)
while the Gibbs picture yields the quantum version of the time-dependent GPG equation (44),
$i\hbar\partial_{t}\hat{\Psi}^{\prime}(\mathbf{x},t)=[\hat{\Psi}^{\prime}(\mathbf{x},t),\hat{K}]=\left[-\frac{\nabla^{2}}{2}+V(\mathbf{x})+\hat{\Psi}^{\prime\dagger}(\mathbf{x},t)\hat{\Psi}^{\prime}(\mathbf{x},t)-\boldsymbol{\lambda}\cdot\mathbf{T}\right]\hat{\Psi}^{\prime}(\mathbf{x},t).$ (67)
Quantum fluctuations around the coherent GPG expectation value can be described by expanding the field operator as $\hat{\Psi}^{\prime}(\mathbf{x},t)=\Psi_{0}(\mathbf{x})+\hat{\varphi}(\mathbf{x},t)$, recovering at linear order the quantum time-dependent BdGG equations (45),
$i\partial_{t}\hat{\Phi}=M_{0}\hat{\Phi},\leavevmode\nobreak\ \hat{\Phi}(\mathbf{x},t)=\left[\begin{array}[]{c}\hat{\varphi}(\mathbf{x},t)\\ \hat{\varphi}^{\dagger}(\mathbf{x},t)\end{array}\right].$ (68)
@@
$\hat{\gamma}_{n},\hat{\gamma}^{\dagger}_{m}]=[(z_{n}|\hat{\Phi}),(\hat{\Phi}|z_{m})]=(z_{n}|z_{m})=\delta_{nm}.$ (71)
@@
$i\partial_{t}\hat{\gamma}_{n}(t)$ $=$ $(z_{n}|M_{0}\hat{\Phi})=\epsilon_{n}\hat{\gamma}_{n}(t)\Longrightarrow$
$\hat{\gamma}_{n}(t)$ $=$ $\hat{\gamma}_{n}e^{-i\epsilon_{n}t}.$ (72)
@@
$\hat{\gamma}_{a},\hat{\gamma}_{b}]=(z_{a}|z_{b})=i\Omega_{ab}\Longrightarrow[\hat{\gamma}^{a},\hat{\gamma}^{b}]=-i\Omega^{ab}.$ (74)
This implies that the components of the vector $\boldsymbol{\hat{\gamma}}=(\hat{\gamma}^{\alpha},\hat{\gamma}^{A})\equiv(\hat{X}^{\alpha},\hat{P}^{A})$ can be seen as generalized coordinate/momentum operators, as those also satisfy the same symplectic algebra:
$\hat{X}^{\alpha},\hat{P}^{B}]=[\hat{X}_{A},\hat{P}^{B}]=i\delta_{A}^{B}.$ (75)
Indeed, we can take this analogy further by analyzing their equations of motion,
[…]
@@
[…]
@@
The quantum fluctuations of the velocities are then translated into a ballistic motion for the quantum amplitude of the NG modes, arriving at Eq. (79). Another way to put it is that the modes $z_{A}$ are not eigenmodes of the BdGG operator $M_{0}$. Instead, from Eq. (23), we have that
@@
which yields $z_{A}(\mathbf{x},t)=z_{A}(\mathbf{x})+(\partial_{A}\lambda^{\beta})z_{\beta}t$ in the time-dependent BdGG equation (45). As a result, we recover the same intuitive picture of the classical mechanical example in Sec. 2: the dynamics of the fluctuations along the orbits generated by the broken symmetry transformations lead to unbounded ballistic motion as there is no restoring force. These results are thus limited to sufficiently short times, when the NG amplitudes $\hat{X}^{\alpha}(t)$ can be treated as small, and the linear BdG approximation is still valid. Nevertheless, our formalism also provides the basis for a full nonperturbative expansion [55, 56], whose study is beyond the scope of the present work.
After gathering Eqs. (3.4), (78), we find that the quantum field fluctuations evolve as
@@
$+\sum_{n}\hat{\gamma}_{n}z_{n}(\mathbf{x})e^{-i\epsilon_{n}t}+\hat{\gamma}^{\dagger}_{n}(t)\bar{z}_{n}(\mathbf{x})e^{i\epsilon_{n}t}.$ (83)
@@
The same dynamics could have been derived from the generalized Gibbs Hamiltonian by inserting Eq. (69) in $\hat{K}=\hat{H}-\boldsymbol{\lambda}\cdot\mathbf{\hat{Q}}$ and expanding up to quadratic order in the field fluctuations, in the spirit of the Bogoliubov approximation. After dropping out all mean-field and zero-point $c$-number contributions, we arrive at
$\hat{K}=\frac{1}{2}(\hat{\Phi}|M_{0}\hat{\Phi})=\frac{\partial^{2}_{AB}E}{2}\hat{P}^{A}\hat{P}^{B}+\sum_{n}\epsilon_{n}\hat{\gamma}^{\dagger}_{n}\hat{\gamma}_{n}.$ (84)
@@
The second term is the usual Bogoliubov contribution while the first one is the Goldstone-Gibbs contribution. The latter is quadratic in the fluctuations of the conserved charges, then behaving as effective momenta, and does not depend on the NG amplitudes, since the Hamiltonian is invariant under the symmetry transformations. In particular, if the solution is energetically stable, then the susceptibility matrix $\partial^{2}_{AB}E$ is a positive definite form. In fact, one can always diagonalize $\partial^{2}_{AB}E$ by performing an appropriate rotation in the Goldstone-Gibbs sector, $\partial^{2}_{AB}E=M_{A}^{-1}\delta_{AB}$, where $M_{A}$ plays the role of an effective mass. This rotation yields a new set of independent Goldstone-Gibbs modes $\{z^{\prime}_{\alpha},z^{\prime}_{A}\}$, whose amplitudes $\{\hat{X}^{\prime\alpha},\hat{P}^{\prime A}\}$ satisfy
@@
$\hat{K}=\frac{(\hat{P}^{\prime A})^{2}}{2M_{A}}+\sum_{n}\epsilon_{n}\hat{\gamma}^{\dagger}_{n}\hat{\gamma}_{n},\leavevmode\nobreak\ \hat{X}^{\prime\alpha}(t)=\hat{X}^{\prime\alpha}+\frac{\hat{P}^{\prime A}}{M_{A}}t.$ (85)
@@
Thus, the susceptibility matrix, computed from thermodynamical considerations, determines how the different NG modes hybridize in a dynamical context. Moreover, as the NG modes are the zero-energy limit of some branches from the regular Bogoliubov sector, the structure of the hybrid Goldstone-Gibbs modes $\{z^{\prime}_{\alpha},z^{\prime}_{A}\}$ also describes the low-energy regime.
@@
The quantization of the Goldstone-Gibbs modes in the usual GP context is simply retrieved by first setting all Lagrange multipliers to zero except for the chemical potential $\lambda^{\theta}=\mu$, so the only independent conserved charge is the particle number $N$. With the resulting GP solution, one computes the remaining charges $Q_{\alpha}$ to assemble the vector $\mathbf{Q}=\mathbf{Q}_{0}$. Once known, the full GPG equation is solved as a function of $\mathbf{Q}$, from where the Berry-Gibbs curvature $F_{AB}$ and the susceptibility matrix $\partial^{2}_{AB}E$ can be locally evaluated at the point $\mathbf{Q}_{0}$ characterizing the original GP solution.
The presence of dynamical instabilities would not modify the main results of this section as these modes are orthogonal to both the Goldstone-Gibbs and the regular Bogoliubov sector. Their quantization is more involved as they behave like unstable oscillators, bearing some resemblance to the Goldstone-Gibbs procedure; see Refs. [78, 79, 80] for more details.
@@
To conclude this section, we stress that the thermodynamical description is essential to ensure the conservation of the commutation rules between the NG amplitudes because it implies that the flow of the GPG wavefunction is potential, Eq. (32), and hence
$\hat{X}_{A}(t),\hat{X}_{B}(t)]=i(\partial_{B}\lambda_{A}-\partial_{A}\lambda_{B})t=0.$ (86)
Figure 1: Density profile $|\Psi_{\mathbf{n}}(x)|^{2}$ of a cnoidal wave for increasing $\nu=0$ (horizontal black), $\nu=0.2$ (solid blue), $\nu=0.7$ (dashed-dotted green), and $\nu=0.9999$ (dashed red), with fixed $J=0.5$ and $\mu_{v}=1.125$.
4 Cnoidal wave in a superfluid